Question:

Suppose that business travelers and vacationers have the following demand for airline tickets from New York to Boston:

PriceQuantity Demanded (business travelers)Quantity Demanded (vacationers)$150 2,100tickets 1,000tickets 200 2,000 800 250 1,900 600 300 1,800 400a. As the price of tickets rises from $200 to $250, what is the price elasticity of demand for (i) business travelers and (ii) vacationers?


1
Expert's answer
2021-10-19T15:39:37-0400



"e_(p) = \frac{dQ}{Q}/\frac{dP}{P}"

Where:

"e_(p)" - price elasticity

"Q" - Quantity f the demanded good

"P" - Price of the demanded good



(I) Price elasticity of demand for business travelers


Price elasticity of demand "=\frac{1,900 -2,000}{1,900+2,000}/\frac {250-200}{250+200}"



Price elasticity of demand "= \frac{-100}{3,900}/\frac{50}{450} = -0.2307"


"e_(p) = 0.2307"


(ii) Price elasticity of demand for vacation travelers


Price elasticity of demand ="=\frac{600-800}{600+800} /\frac{250-200}{250+200}"



Price elasticity of demand ="= \frac{-200}{1400}/\frac{50}{450} = -1.2867"


"e_(p) =1.2867"